Convert from IEE 754 32-bit format to decimal

a) 0100001010100000...0

b) 10111111010000...0

Convert from decimal to IEE 754 32bit format

a) -(7/16)
b) 137
Answer #1

// I hope this is your requirement

//Dont dislike

Convert from IEEE 754 32-bit format to decimal

a) 0100001010100000...0= 80

b) 10111111010000...0= -0.75

Convert from decimal to IEEE 754 32-bit format

a) -(7/16)= 1-01111101-11000000000000000000000
b) 137= 0-10000110-00010010000000000000000